# A Cost-Efficient FPGA Implementation of Tiny Transformer Model using Neural ODE

### v2

@@ Abstract
Transformer has been adopted to a wide range of tasks and shown to outperform CNNs and RNNs while it suffers from high training cost and computational complexity. To address these issues, a hybrid approach has become a recent research trend, which replaces a part of ResNet with an MHSA (Multi-Head Self-Attention). In this paper, we propose a lightweight hybrid model which uses Neural ODE (Ordinary Differential Equation) as a backbone instead of ResNet for 12.1$\times$ parameter reduction. For the STL10 dataset, the proposed model achieves 80.15% top-1 accuracy which is comparable to ResNet50. Then, the proposed model is deployed on a modest-sized FPGA device for edge computing. To further reduce FPGA resource utilization, the model is quantized following QAT (Quantization Aware Training) scheme instead of PTQ (Post Training Quantization) to suppress the accuracy loss. As a result, an extremely lightweight Transformer-based model can be implemented on resource-limited FPGAs. The weights of the feature extraction network are stored on-chip to minimize the memory transfer overhead, allowing faster inference. By eliminating the overhead of memory transfers, inference can be executed seamlessly, leading to accelerated inference. The proposed FPGA implementation achieves a 34.01$\times$ speedup for the backbone and MHSA parts, and it achieves an overall 9.85$\times$ speedup when taking into account software pre- and post-processing. It also achieves an overall 7.10$\times$ higher energy efficiency compared to the ARM Cortex-A53 CPU.
@@ 1 Introduction
Among these models, we focus on BoTNet (Bottleneck Transformer Network) [8], which replaces spatial convolutions in the last three bottleneck blocks in ResNet [11] with the global self-attention and outperforms baseline models by a large margin. It achieves 84.7% top-1 accuracy on ImageNet benchmark.
Based on the growing number of IoT devices, the demand for deploying DNNs to edge devices (e.g., self-driving cars, security cameras, and drones) has increased. FPGA is becoming an attractive platform to accelerate DNNs due to its energy efficiency, reconfigurability, and short development period. We aim to implement the proposed model on a modest-sized FPGA such as Xilinx ZCU104. In our previous work [12], MHSA mechanism was implemented on the programmable logic part of the FPGA; since a large part of the model was still performed on the host CPU, only a moderate performance improvement was achieved. In contrast, this paper implements the entire feature extraction network of the proposed model on the FPGA to fully harness its computational capability. Additionally, to reduce FPGA resources, in this paper we choose to employ a lightweight quantization technique by considering the characteristics of Neural ODE. The parameter reduction and model quantization play an important role to save the scarce on-chip memory and map the Transformer-based model onto such FPGAs.
In this paper, we propose a tiny Transformer-based model (see Fig. 1) for resource-constrained edge devices by approximating the ResNet computation with Neural ODE [13], which reformulates the forward propagation of ResNet blocks as a numerical solution of ODEs. The number of parameters is greatly reduced because the forward pass through multiple ResNet building blocks is replaced by an iteration of the same single block, as shown in Fig. 1 (bottom).
Figure 1: The proposed Transformer-based model on FPGA
@@ 2.1 CNN and Transformer
In recent years, numerous Transformer-based models have been proposed for image recognition tasks. ViT [2] is a pure Transformer-based model for image recognition and achieves greater accuracy than CNN-based models with a large dataset (e.g., JFT-300M). However, one major drawback of ViT is that it performs worse than CNN-based models when trained with small- to medium-scale datasets such as CIFAR-10/100 [14] and ImageNet [15] due to the lack of intrinsic inductive biases.
@@ 2.2 Reduction of model computational complexity
Several extensions to Neural ODE have been proposed in the literature. Neural SDE (Stochastic Differential Equation) [24] incorporates various common regularization mechanisms based on the introduction of random noise. This model shows greater robustness to input noise, whether they are intentionally adversarial or not, compared to Neural ODE. In [25], training cost is significantly reduced without sacrificing accuracy by applying two appropriate regularizations to Neural ODE. In [26], by considering the regularization, TisODE (Time-invariant steady Neural ODE) is proposed to further enhance the robustness of vanilla Neural ODE. In [27], inherent constraints in Neural ODE are identified, and more expressive model called ANODE (Augmented Neural ODE) is proposed. These approaches are crucial in achieving robustness, approximation capabilities, and performance improvements in the context of Neural ODE. While this paper incorporates Transformer architecture into a model with vanilla Neural ODE as a backbone, the same idea can be applied to these Neural ODE extensions as well, which is not the focus of this paper. To further enhance the proposed model performance, the above techniques are incorporated as future work.
@@ 2.4 Research gap and objectives
In the midst of the research trend on Transformer and CNN hybrid models [5, 6, 7, 8, 9, 10, 16, 17, 18], there are few proposals for small-scale hybrid models aimed at edge computing. The objective of this paper is to demonstrate a lightweight hybrid model of CNN and attention mechanism which can be implemented on resource-limited FPGA devices.
@@
In [12], a lightweight Transformer model with Neural ODE is proposed along with its FPGA-based design. Since only an MHSA block is accelerated and the other part including Neural ODE blocks is still executed on a host CPU, the design shows only a limited performance gain. Contrary to that, in this paper, we implement the entire feature extraction network to further improve the inference speed and fully benefit from the computing power of FPGAs. Additionally, in [12], the model is quantized by simply converting inputs/outputs and weights from floating- to fixed-point representation. However, it is shown that performing such simple type conversions throughout the entire model would result in a significant degradation of accuracy [12]. In this paper, QAT is instead employed, which allows more aggressive bit-width reduction (e.g., 8-bit quantization) to save the on-chip memory while mitigating the risk of significant accuracy loss.
@@ 3.1.3 Positional encoding
Since the self-attention (Eq. (7)) is a set operation and is invariant to the random permutation of inputs, it causes a loss of the positional information. For example, in Eq. (2), if $\mathbf{x}_{i}$ and $\mathbf{x}_{j}$ are swapped, then their corresponding output elements, i.e., $\mathbf{a}_{i}$ and $\mathbf{a}_{j}$, are simply swapped. The model with this property, called equivariance, is not appropriate for structural data such as images, since the model should be aware of the spatial relation of the pixels. To address this information loss, a positional encoding is often employed with the attention mechanism. A position-specific vector $\mathbf{p}_{i}\in\mathbb{R}^{D}$ is added to the $i$-th input $\mathbf{x}_{i}\in\mathbb{R}^{D}$ via addition $\mathbf{x}_{i}+\mathbf{p}_{i}$ or concatenation $\left[\mathbf{x}_{i},\mathbf{p}_{i}\right]$. The positional encoding is modeled as a deterministic function with a set of fixed hyperparameters or as a learnable function. In addition, either absolute or relative position encoding can be used. Typically, Transformer [1] adopts the sinusoidal function, which is categorized as a fixed absolute encoding and is written as follows:
@@
where $j=1,\dots,D/2$. On the other hand, several studies [2, 23] propose to learn the relative positional encoding via MLP. In [8], the authors show that the relative position leads to better accuracy than the absolute one. Our method thus uses a learnable relative encoding.
@@ 3.1.4 MHSA
$\begin{split}\mathrm{MHSA}\mathbf{(X)}=[SA_{1}(\mathbf{X});SA_{2}(\mathbf{X});...;SA_{k}(\mathbf{X})]\in\mathbb{R}^{N\times kD_{h}}\end{split}$ (12)
@@ 3.3 Quantization of the proposed model
While there are a number of quantization methods, we employ a method based on a learnable lookup table (LLT) [32], considering its simplicity and low computational overhead. LLT-based quantization is outlined in the following.
@@
In the following, we describe the quantization process for activations. Note that weights can be quantized in the same way. The full-precision activation $\hat{a}\in[0,1]$ is converted to its nearest discrete value $i/2^{n}$, which is represented by a total of $12^{n}-$ 1 learnable step functions. Here, when $\bar{a}$ is within the range $[i/2^{n},(i+1)/2^{n}]$, it is mapped to either $i/2^{n}$ or $(i+1)/2^{n}$ by the $i$-th step function $\mathrm{step}_{i}(.)$ ($i=0,1,\ldots,2^{n}-1$) as follows:
@@ 4 Design of the proposed model
(1) ResNet consists of pre-processing, post-processing, and multiple stages for feature extraction as shown in Table 1. Pre-processing converts input images into feature maps, and post-processing transforms feature maps into a vector with elements equal to the number of classes. Each stage comprises several ResBlocks with the same structure. Specifically, each ResBlock contains two sets of convolution (Conv) layer, batch normalization (BatchNorm) layer, and activation function (e.g., ReLU), which are cascaded sequentially. In each stage, the first ResBlock is employed for downsampling, denoted as “DS”, followed by a stack of standard ResBlocks. The ResBlock for DS halves the width and height of a feature map while doubling the number of channels, i.e., it takes the feature map of size $(Ch,H,W)$ as input and produces an output of size $(2Ch,H/2,W/2)$. Note that the ResBlock for DS and standard ResBlock only differ in the configuration of the first convolution layer. For example, a straightforward approach to achieve DS is to set the stride and output channels of the first convolution in the block to 2 and $2Ch$, respectively.
@@
Another modification is to use ReLU [^1] As an activation function, ReLU6 is also tested, and it shows a comparable accuracy to ReLU. In this paper, ReLU is used as well as [37]. Using ReLU6 for higher robustness is our future work. instead of softmax as an activation function. As described in Sec. 3.1, relationships between query and key (i.e., logits) are obtained by taking an inner-product of $\mathbf{Q}$ with $\mathbf{K}$, and then they are passed on to the softmax function such that the attention weights $a_{i,1},\ldots,a_{i,N}$ for the $i$-th query sum to one. An immediate advantage of this modification is that ReLU is hardware-friendly as it only consumes one comparator and one multiplexer. According to [37], the accuracy of ReLU-based attention is comparable to that of the original softmax-based one. In addition, the attention weights become sparser, which assists the analysis of information flow in the model. Besides, LayerNorm [38] is added at the end of MHSABlock to stabilize gradients and facilitate the convergence.
@@
$\begin{split}\mathrm{MHSA(\mathbf{X})}=\mathrm{LayerNorm}\left([SA_{1}(\mathbf{X});SA_{2}(\mathbf{X});...;SA_{k}(\mathbf{X})]\right).\end{split}$ (29)
@@ 5 Implementation
As described in Sec. 4, the proposed model significantly reduces parameter size, and thus essential parameters for the feature extraction part can be stored in on-chip BRAM and URAM buffers. That is, only input and output data are transferred between the PL part and DRAM during inference. In this case, because the data transfer amount is quite small, the performance of the proposed design is compute-bound. In the FPGA implementation, the model assumes the input image resolution of 96 $\times$ 96. Note that the input image size has less impact on the model’s parameter size because the feature map size does not affect the parameter size of convolution layers.
@@ 5.1 Board-level implementation
Fig. 5 depicts the block diagram of the proposed software-hardware co-design. The feature extraction part of the proposed model (i.e., from ODEBlock1 to MHSABlock) is implemented as a custom IP core. The custom IP core along with a DMA (Direct Memory Access) controller is implemented on the PL part. The DMA controller (DMAC) is in charge of transferring the input/output feature maps as well as the model parameters (e.g., LUTs for activations, quantized weights, and relative positional encoding) via AXI-Stream, and is connected to the 128-bit High-Performance (HP0) port. The IP core provides control registers for setting necessary configuration such as the number of ODE iterations. The PS part accesses these control registers through a 32-bit AXI-Lite interface (Fig. 5, blue lines). The pre- and post-processing blocks are also executed on the PS part.
@@ 5.2 FPGA implementation of the proposed model
The proposed model consists of three types of blocks: ODEBlock, DSBlock, and MHSABlock. The proposed model is designed with C/C++ with Vitis HLS. The proposed accelerator is implemented as a stream of blocks, as shown in Fig. 6. That is, input samples go through separated blocks, such as ODEBlock1, DSBlock1, ODEBlock2, DSBlock2, and MHSABlock. All layers in each block are parallelized along the output channel dimension by utilizing multiple DSP blocks. We describe the implementation of ODEBlock, DSBlock, and MHSABlock in the following.

### v3

# A Cost-Efficient FPGA-Based CNN-Transformer using Neural ODE

@@ Abstract
Transformer has been adopted to image recognition tasks and shown to outperform CNNs and RNNs while it suffers from high training cost and computational complexity. To address these issues, a hybrid approach has become a recent research trend, which replaces a part of ResNet with an MHSA (Multi-Head Self-Attention). In this paper, we propose a lightweight hybrid model which uses Neural ODE (Ordinary Differential Equation) as a backbone instead of ResNet so that we can increase the number of iterations of building blocks while reusing the same parameters, mitigating the increase in parameter size per iteration. The proposed model is deployed on a modest-sized FPGA device for edge computing. The model is further quantized by QAT (Quantization Aware Training) scheme to reduce FPGA resource utilization while suppressing the accuracy loss. The quantized model achieves 79.68% top-1 accuracy for STL10 dataset that contains 96$\times$96 pixel images. The weights of the feature extraction network are stored on-chip to minimize the memory transfer overhead, allowing faster inference. By eliminating the overhead of memory transfers, inference can be executed seamlessly, leading to accelerated inference. The proposed FPGA implementation accelerates the backbone and MHSA parts by 34.01$\times$, and achieves an overall 9.85$\times$ speedup when taking into account the software pre- and post-processing. The FPGA acceleration leads to 7.10$\times$ better energy efficiency compared to the ARM Cortex-A53 CPU. The proposed lightweight Transformer model is demonstrated on Xilinx ZCU104 board for the image recognition of 96$\times$96 pixel images in this paper and can be applied to different image sizes by modifying the pre-processing layer.
@@ 1 Introduction
Among these models, BoTNet (Bottleneck Transformer Network) [8] replaces spatial convolutions in the last three bottleneck blocks in ResNet [11] with the global self-attention (i.e., MHSA) and outperforms baseline models by a large margin. We focus on BoTNet because it has shown an improved accuracy in object detection, instance segmentation, and image classification, demonstrating the benefits of including MHSA blocks in the CNN.
As mentioned above, the Transformer-based models are attractive in terms of accuracy while the computation cost is high (e.g., computation cost of MHSA grows with the square of input size $N$). FPGA is becoming an attractive platform to accelerate DNNs due to its energy efficiency, reconfigurability, and short development period. We thus aim to implement the proposed model including MHSA on a modest-sized FPGA such as Xilinx ZCU104. In this paper, we propose a tiny Transformer-based model (see Fig. 1) for resource-constrained FPGA devices by approximating the ResNet computation with Neural ODE [12], which reformulates the forward propagation of ResNet blocks as a numerical solution of ODEs. The number of parameters is greatly reduced because the forward pass through multiple ResNet building blocks is replaced by an iteration of the same single block, as shown in Fig. 1 (bottom). We can thus implement a number of iterations of building blocks while reusing the same parameters, significantly reducing the parameter size compared to ResNets.
In our previous work [13], MHSA mechanism was implemented on the programmable logic part of the FPGA; since a large part of the model was still performed on the host CPU, only a moderate performance improvement was achieved. In contrast, this paper implements the entire feature extraction network of the proposed model on the FPGA to fully harness its computational capability. Additionally, in this paper we choose to employ a lightweight quantization technique by considering the characteristics of Neural ODE, because the parameter reduction and model quantization play an important role to save the scarce on-chip memory and map the Transformer-based model onto such FPGAs. Specifically, the model is quantized by QAT scheme instead of PTQ (Post Training Quantization) to further reduce FPGA resources while suppressing the accuracy loss.
Figure 1: Overview of the proposed system
@@ 2.1 CNN and Transformer
CNNs have been used for various image recognition tasks. For example, various CNN models are tested for detecting COVID-19 infection [14, 15]. In addition, numerous Transformer-based models have been recently proposed for image recognition tasks. ViT [2] is a pure Transformer-based model for image recognition and achieves greater accuracy than CNN-based models with a large dataset (e.g., JFT-300M). However, one major drawback of ViT is that it performs worse than CNN-based models when trained with small- to medium-scale datasets such as CIFAR-10/100 [16] and ImageNet [17] due to the lack of intrinsic inductive biases.
@@ 2.2 Reduction of model computational complexity
Several extensions to Neural ODE have been proposed in the literature. Neural SDE (Stochastic Differential Equation) [26] incorporates various common regularization mechanisms based on the introduction of random noise. This model shows greater robustness to input noise, whether they are intentionally adversarial or not, compared to Neural ODE. In [27], training cost is significantly reduced without sacrificing accuracy by applying two appropriate regularizations to Neural ODE. In [28], by considering the regularization, TisODE (Time-invariant steady Neural ODE) is proposed to further enhance the robustness of vanilla Neural ODE. In [29], inherent constraints in Neural ODE are identified, and more expressive model called ANODE (Augmented Neural ODE) is proposed. These approaches are crucial in achieving robustness, approximation capabilities, and performance improvements in the context of Neural ODE. While this paper incorporates Transformer architecture into a model with vanilla Neural ODE as a backbone, the same idea can be applied to these Neural ODE extensions as well, though such exploration is beyond the scope of this paper.
@@ 2.4 Research gap and objectives
In the midst of the research trend on Transformer and CNN hybrid models [5, 6, 7, 8, 9, 10, 18, 19, 20], there are few proposals for small-scale hybrid models aimed at edge computing. Although tiny Transformer models [34, 35] have been proposed for edge devices, they are mostly for signal processing and not for image recognition tasks. The objective of this paper is thus to demonstrate a lightweight hybrid model of CNN and attention mechanism which can be implemented on resource-limited FPGA devices. To fill in the research gap toward the objective, this paper introduces the Neural ODE method for the hybrid model and demonstrates the benefits of our approach on the modest-sized FPGA device for edge computing.
@@
2.5 Implementation challenges
In [13], a lightweight Transformer model with Neural ODE is proposed along with its FPGA-based design. Since only an MHSA block is accelerated and the other part including Neural ODE blocks is still executed on a host CPU in [13], the design shows only a limited performance gain. Contrary to that, in this paper, our new approach is to implement the entire feature extraction network to further optimize the inference speed and fully benefit from the computing power of FPGAs. Another implementation challenge is that, although the model in [13] is quantized by simply converting inputs/outputs and weights from floating- to fixed-point representation, such simple type conversions throughout the entire model would result in a significant degradation of accuracy [13]. To address this issue, in this paper, QAT is instead employed, which allows more aggressive bit-width reduction (e.g., 8-bit quantization) to save the on-chip memory while mitigating the risk of significant accuracy loss.
@@ 3.1.3 Positional encoding
Since the self-attention (Eq. (7)) is a set operation and is invariant to the random permutation of inputs, it causes a loss of the positional information. For example, in Eq. (2), if $\mathbf{x}_{i}$ and $\mathbf{x}_{j}$ are swapped, then their corresponding output elements, i.e., $\mathbf{a}_{i}$ and $\mathbf{a}_{j}$, are simply swapped. The model with this property, called equivariance [2], is not appropriate for structural data such as images, since the model should be aware of the spatial relation of the pixels. To address this information loss, a positional encoding is often employed with the attention mechanism. A position-specific vector $\mathbf{p}_{i}\in\mathbb{R}^{D}$ is added to the $i$-th input $\mathbf{x}_{i}\in\mathbb{R}^{D}$ via addition $\mathbf{x}_{i}+\mathbf{p}_{i}$ or concatenation $\left[\mathbf{x}_{i},\mathbf{p}_{i}\right]$. The positional encoding is modeled as a deterministic function with a set of fixed hyperparameters or as a learnable function. In addition, either absolute or relative position encoding can be used. Typically, Transformer [1] adopts the sinusoidal function, which is categorized as a fixed absolute encoding and is written as follows:
@@
where $j=1,\dots,D/2$. On the other hand, several studies [2, 25] propose to learn the relative positional encoding via MLP. In [8], the authors show that the relative position leads to better accuracy than the absolute one. Thus, a learnable relative encoding is used in this study.
@@ 3.1.4 MHSA
$\begin{split}\mathrm{MHSA}\mathbf{(X)}=[SA_{1}(\mathbf{X});...;SA_{k}(\mathbf{X})]\in\mathbb{R}^{N\times kD_{h}}\end{split}$ (12)
@@ 3.3 Quantization of the proposed model
While there are a number of quantization methods, in this paper a learnable lookup table (LLT) based method [36] is employed, considering its simplicity and low computational overhead. LLT-based quantization is outlined in the following.
@@
In the following, the quantization process for activations is described. Note that weights can be quantized in the same way. The full-precision activation $\hat{a}\in[0,1]$ is converted to its nearest discrete value $i/2^{n}$, which is represented by a total of $12^{n}-$ 1 learnable step functions. Here, when $\bar{a}$ is within the range $[i/2^{n},(i+1)/2^{n}]$, it is mapped to either $i/2^{n}$ or $(i+1)/2^{n}$ by the $i$-th step function $\mathrm{step}_{i}(.)$ ($i=0,1,\ldots,2^{n}-1$) as follows:
@@ 4 Design of the proposed model
(1) ResNet consists of pre-processing, post-processing, and multiple stages for feature extraction as shown in Table 1. Pre-processing converts input images into feature maps, and post-processing transforms feature maps into a vector with elements equal to the number of classes. Each stage comprises several ResBlocks with the same structure. Specifically, each ResBlock contains two sets of convolution (Conv) layer, batch normalization (BatchNorm) layer, and activation function (e.g., ReLU), which are cascaded sequentially. In each stage, the first ResBlock is employed for downsampling, denoted as “DS”, followed by a stack of standard ResBlocks. The ResBlock for DS halves the width and height of a feature map while doubling the number of channels, i.e., it takes the feature map of size $(Ch,H,W)$ as input and produces an output of size $(2Ch,H/2,W/2)$  [^1] It is expected that the DS block reduces the feature map size while extracting more complex features by increasing the number of channels as the layer goes deeper. This structure is commonly used in well-known CNNs such as VGG [37], ResNet [11], and U-Net [38]. We thus follow this approach in our proposed model.. Note that the ResBlock for DS and standard ResBlock only differ in the configuration of the first convolution layer. For example, a straightforward approach to achieve DS is to set the stride and output channels of the first convolution in the block to 2 and $2Ch$, respectively.
@@
Another modification is to use ReLU [^2] As an activation function, ReLU6 is also tested, and it shows a comparable accuracy to ReLU. In this paper, ReLU is used as well as [43]. instead of softmax as an activation function. As described in Sec. 3.1, relationships between query and key (i.e., logits) are obtained by taking an inner-product of $\mathbf{Q}$ with $\mathbf{K}$, and then they are passed on to the softmax function such that the attention weights $a_{i,1},\ldots,a_{i,N}$ for the $i$-th query sum to one. An immediate advantage of this modification is that ReLU is hardware-friendly as it only consumes one comparator and one multiplexer. According to [43], the accuracy of ReLU-based attention is comparable to that of the original softmax-based one. In addition, the attention weights become sparser, which assists the analysis of information flow in the model. Besides, LayerNorm [44] is added at the end of MHSABlock to stabilize gradients and facilitate the convergence.
@@
$\begin{split}\mathrm{MHSA(\mathbf{X})}=\mathrm{LayerNorm}\left([SA_{1}(\mathbf{X});...;SA_{k}(\mathbf{X})]\right).\end{split}$ (29)
@@ 5 Implementation
As described in Sec. 4, the proposed model significantly reduces parameter size, and thus essential parameters for the feature extraction part can be stored in on-chip BRAM and URAM buffers. Only input and output data are transferred between the PL part and DRAM during inference. In this case, because the data transfer amount is quite small, the performance of the proposed design is compute-bound. In the FPGA implementation, the model assumes the input image resolution of 96 $\times$ 96. Note that the input image size has less impact on the model’s parameter size because the feature map size does not affect the parameter size of convolution layers.
@@ 5.1 Board-level implementation
Fig. 5 depicts the block diagram of the proposed software-hardware co-design. The feature extraction part of the proposed model (i.e., from ODEBlock1 to MHSABlock) is designed as a custom IP core and implemented on the PL part. The IP core transfers input/output feature maps as well as model parameters (e.g., LUTs for activations, quantized weights, and relative positional encoding) via the AXI master interface connected to the 128-bit High-Performance (HP0) port. It also provides control registers for setting necessary configuration parameters such as the number of ODE iterations. The PS part accesses these control registers through a 32-bit AXI-Lite interface (Fig. 5, blue lines). The pre- and post-processing blocks are also executed on the PS part.
@@ 5.2 FPGA implementation of the proposed model
The proposed model consists of three types of blocks: ODEBlock, DSBlock, and MHSABlock. The proposed model is developed in C++ with Vitis HLS 2023.1 and implemented with Vivado 2023.1. The proposed accelerator consists of a stream of blocks, as shown in Fig. 6. Input samples are processed by separated blocks in order, namely ODEBlock1, DSBlock1, ODEBlock2, DSBlock2, and MHSABlock. These blocks use shared buffers for the communication between them. Specifically, ODEBlock1 reads input features from the first buffer, computes output features, and stores them to the next buffer. It repeats these steps until all the input features are processed. Then, DSBlock1 reads input features from the second buffer and computes output features as in the previous block. These shared buffers are implemented with BRAMs and reused as much as possible to save the on-chip BRAM resources. All layers in each block are parallelized along the output channel dimension by utilizing multiple DSP blocks. We describe the implementation of ODEBlock, DSBlock, and MHSABlock in the following.